### v1

# Click Without Compromise: Online Advertising Measurement via Per User Differential Privacy

Abstract.
Online advertising is a cornerstone of the Internet ecosystem, with advertising measurement playing a crucial role in optimizing efficiency. Ad measurement entails attributing desired behaviors, such as purchases, to ad exposures across various platforms, necessitating the collection of user activities across these platforms. As this practice faces increasing restrictions due to rising privacy concerns, safeguarding user privacy in this context is imperative. Our work is the first to formulate the real-world challenge of advertising measurement systems with real-time reporting of streaming data in advertising campaigns. We introduce Ads-BPC, a novel user-level differential privacy protection scheme for advertising measurement results. This approach optimizes global noise power and results in a non-identically distributed noise distribution that preserves differential privacy while enhancing measurement accuracy. Through experiments on both real-world advertising campaigns and synthetic datasets, Ads-BPC achieves a 25% to 50% increase in accuracy over existing streaming DP mechanisms applied to advertising measurement. This highlights our method’s effectiveness in achieving superior accuracy alongside a formal privacy guarantee, thereby advancing the state-of-the-art in privacy-preserving advertising measurement.
Differential Privacy, User Level Privacy, Advertising Measurement, Impression, Conversion, Multi-touch Attribution, Online Streaming Dataset
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
1. Introduction
Online advertising is essential in the Internet ecosystem because it fuels the economy of the web and enables businesses to connect with their target audience effectively. Advertising measurement systems (yun2020challenges, ; guha2010challenges, ; briggs1997advertising, ; yuan2013real, ) are pivotal in the online advertising landscape. They aim to assess the effectiveness of advertising campaigns by determining the total number of conversions—such as purchases, sign-ups, downloads, or other desired actions—that occur as a direct result of ad exposure. To reach a more diverse audience, advertisers often set up ad campaigns across various platforms, including Google Ads, Facebook Ads, and other digital advertising networks. To effectively monitor the impact of these ads on each platform, appropriate tracking mechanisms are in place to allow for real-time monitoring of advertising campaigns across the selected channels.
The ads measurement process encompasses two major steps: attribution and computation. The attribution step involves analyzing the customer’s online journey, starting from initial ad exposure to the ultimate conversion. Attribution aims to identify which exposures (touchpoints) contributed to conversions. The method that allocates credit to various advertising channels or touchpoints throughout the customer journey is referred to as multi-touch attribution (MTA). MTA recognizes the multiple interactions a customer may have with different advertising channels before making a conversion. By distributing credit among these touchpoints based on their influence on the conversion, MTA models offer a more nuanced understanding of each channel’s effectiveness within the advertising mix. The computation step is to calculate the total conversion value based on the attribution results to gain insights into the effectiveness of different advertising channels. Based on the such insights, advertisers can improve their advertising campaigns by reallocating budget, adjusting targeting parameters, refining ad designs, or tweaking other campaign elements to improve their return on investment and user advertising experience.
Ad measurement requires collecting a variety of user data, including the ads users have viewed, clicks registered by the ad platform, and conversion data from the advertiser’s end. Traditionally, this involves automatically tracking user behavior across both ad publishers’ platforms and advertisers’ platforms, utilizing tools such as pixels, tags, or software development kits (SDKs). However, the tracking and collection of user data across sites have raised increasing privacy concerns regarding user consent, data security, and the potential misuse of personal information (toubiana2010adnostic, ; yun2020challenges, ; haddadi2011targeted, ; ullah2020privacy, ). These concerns have led to legislative actions, such as the General Data Protection Regulation (GDPR) in Europe (voigt2017eu, ) and industry initiatives such as Apple’s App Tracking Transparency (AppleATT, ), which impose varying degrees of restrictions on cross-site tracking. Recent studies have developed privacy-preserving mechanisms to support advertising measurement (ayala2022show, ; pfeiffer2021masked, ; rosales2012post, ). However, these approaches lack formal privacy guarantees. Our work addresses this gap by introducing a formal differential privacy guarantee for advertising measurement outcomes.
Differential privacy (dwork06Calibrating, ; dworkKMM06:ourdata, ; zcdp, ; renyidp, ; fdp, ) provides a mathematical framework for adding carefully calibrated noise to query results or datasets to protect individual data record while still allowing for meaningful data analysis. The calibrated noise present makes it nearly impossible to discern whether the dataset contains a specific data record, thereby safeguarding sensitive information. Differential privacy has become a key component in modern privacy protection efforts in research areas such as outsourced database systems (near2021differential, ), machine learning model training data protection (abadi2016deep, ), healthcare (dankar2013practicing, ), census statistics (dwork2010differential, ; abowd2018uscensus, ), and data collection (erlingsson2014rappor, ; feldman2023stronger, ).
In this paper, we undertake the novel task of designing a differentially private mechanism to protect individual user privacy in MTA ad measurement systems. Employing differential privacy in such practical settings presents two challenges. First, maintaining a continuous data stream poses significant hurdles for the effective implementation of differential privacy. In advertising campaigns, measurement results must be continuously disclosed to both advertisers and ad providers in real-time, per time slot (e.g., daily or hourly). This is crucial as it allows advertisers to observe the outcomes of their campaigns and adjust their strategies accordingly. Unlike one-time data releases in database queries, continuous streams gradually consume the differential privacy budget, leading to significant accuracy deterioration and making the analysis results less valuable from a business standpoint.
Second, MTA involves records from various platforms, making traditional event-level differential privacy insufficient. In differential privacy, the granularity of privacy refers to the level at which DP guarantees are provided, indicating the unit of data protected by DP. Event-level DP protects whether a specific user action (such as clicking an ad or making a purchase) is included in the dataset, whereas user-level DP provides privacy protection at the level of individual users, covering all event-level records such as views, clicks, and conversions of a particular user. Consequently, user-level DP protects whether a user engaged with the ad campaign at any point, offering a stronger and more meaningful level of privacy protection than event-level DP. For instance, protecting whether a user has ever engaged with an ad for diabetic medication is more meaningful than merely protecting whether a user has clicked on an Instagram ad for diabetic medication.
In our pursuit of providing stronger DP protection for users in practical online advertising systems, this paper presents the following contributions:
  1. (1)
    We are the first to introduce and formalize the problem regarding MTA advertising measurement systems, including both attribution and computation steps, which require the release of measurement results on a streaming basis.
  2. (2)
    We devised a user-level streaming differentially private algorithm, Ads-BPC, tailored to address the challenges in MTA advertising measurement. Mathematically, we formulate the noise calibration process as a total noise variance minimization problem and derive a closed-form solution that adheres to a predetermined privacy budget. This non-identically distributed noise distribution preserves differential privacy and increases measurement accuracy.
  3. (3)
    Through comprehensive experiments with both synthetic and real-world datasets, our algorithm achieves notable improvements in accuracy, ranging from 25% to 50%, compared to previous works. Our results indicate that our algorithm can provide more accurate and meaningful insights for advertisers while ensuring users’ privacy.
This paper is structured as follows: Section 2 illustrates how advertising measurement operates through concrete examples and discuss related concepts in differential privacy. Section 3 explores previous work in the differential privacy area. Section 4 formally defines our user-level differentially private advertising measurement system, formulates it as a convex optimization problem, and presents an efficient algorithm to address it. Section 5 presents discussions on the proposed method. Finally, Section 6 presents the results of comprehensive experiments conducted on both synthetic and real-world datasets using our algorithm.
2. Ads Measurement and Differential Privacy
In this section, we elucidate advertising measurement with multi-touch attribution as well as differential privacy definitions in the advertising measurement context. Mathematical notations used throughout this paper are summarized in Table 1.
Table 1. Table of Notation
$n$: Number of days.
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$: Data vector representation of a stream.
$\mathcal{N}$: All pairs of neighboring streams. ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}}}$: All possible deltas between neighboring. ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$.
$\mathbf{d}$: The difference between two neighboring dataset, $\mathbf{d}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}-\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$.
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u,k}$: $k$-th impression of user $u$.
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u,k}$: $k$-th conversion of user $u$.
$w$: Attribution weight.
$\mathbf{W}$: Prefix-sum matrix.
$\mathbf{L}$: Linear transformation matrix.
$\mathbf{B}$: Basis matrix (strategy matrix).
$\mathbf{z}$: Random variable drawn from a Gaussian distribution.
$\rho$: Privacy budget.
2.1. An Example of Advertising Measurement
At the core of advertising measurement lies the concept of impression and conversion.
Impression Events. An impression event encompasses several key elements that define the context of an ad being served to a user. These elements include: 1) User ID, 2) Publisher ID, 3) Advertiser ID, 4) Timestamp, and 5) Metadata (supplementary information related to the impression, such as the type of user interaction (click or view) and ad format, providing the context for further analysis). We use the notation ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u,k}$ to express the $k$-th impression of user $u$, and use ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\phi}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u,k})$ to express the metadata of the impression ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u,k}$. In the case where an advertiser has multiple products, we use different Advertiser ID for different products. Table 2 shows an example of the impression table, where each row represents an impression event of users interacting (viewing or clicking) with ads on various publishers.
Table 2. Impression Table I on Day 1
User-ID Pub-ID Ad-ID Imp. Time Metadata
$u_{1}$ P-1 Ad-1 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{1},1}$ 1 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\phi}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{1},1})$
$u_{1}$ P-1 Ad-1 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{1},2}$ 11 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\phi}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{1},2})$
$u_{2}$ P-1 Ad-1 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ 15 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\phi}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1})$
$u_{2}$ P-2 Ad-1 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},2}$ 25 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\phi}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},2})$
Conversion Events. Conversion events represent pivotal moments in the advertising pipeline when users generate value for advertisers after interacting with ads on the publishers’ platforms. These events encompass 1) User ID, 2) Advertiser ID, 3) Timestamp, and 4) Metadata: Additional data that characterize the conversion, including its type (e.g., purchase or sign-up) and value (e.g., purchase amount). We use the notation ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u,k}$ to express the $k$-th conversion of user $u$, and use ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\phi}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u,k})$ to express the metadata of the conversion ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u,k}$. Table 3 shows an example of the conversion table, where each row represents a conversion, such as a purchase of Nike shoes.
Table 3. Conversion Table C on Day 1
User-ID Ad-ID Conversion Time Metadata
$u_{1}$ Ad-1 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{1},1}$ 10 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\phi}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{1},1})$
$u_{2}$ Ad-1 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{2},1}$ 20 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\phi}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{2},1})$
$u_{2}$ Ad-1 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{2},2}$ 30 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\phi}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{2},2})$
Advertising Measurement Process The gist of ad measurement can be distilled into a two-step process. In the first step, conversions are attributed to one or more impressions using predefined logic or attribution models. This step is referred to as “attribution,” determining which impressions played a role in driving the eventual conversion event. Section 2.1.1 gives more details about attribution logic. Subsequently, in the second step, queries are executed on the resulting attributed dataset to extract actionable insights. These queries may involve counting the total number of conversions attributed to a subset of impressions, calculating the corresponding conversion rate, or assessing the total return on ad spend for that particular subset.
Table 4. Table after Joining Impression Table I and Conversion Table C for Day 1 based on User-ID and Ad-ID
User-ID Pub-ID Ad-ID Imp. Conv. I.Time C.Time
$u_{1}$ P-1 Ad-1 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{1},1}$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{1},1}$ 1 10
$u_{2}$ P-1 Ad-1 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{2},1}$ 15 20
$u_{2}$ P-1 Ad-1 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{2},2}$ 15 30
$u_{2}$ P-2 Ad-1 ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},2}$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{2},2}$ 25 30
2.1.1. Attribution Logic
For attribution, we assume there is a trusted third party that joins the impression table and the conversion table. Listing 1 shows a simplified SQL query that joins these two tables. Table 4 shows the result of joining Tables 2 and 3. Then the third party assigns the attribution weight of a conversion to each relevant impression. The following are several commonly used attribution logic.
1SELECT User-ID, Pub-ID, Ad-ID,
2Impression, Conversion
3FROM I JOIN C
4ON I.User-ID = C.User-ID
5AND I.Ad-ID = C.Ad-ID
6And I.Time < C.Time
Listing 1: SQL Query to Join Impression Table and Conversion Table on Day 1
Last-Touch Attribution (LTA), also known as Last-Click Attribution, assigns full credit for a conversion to the last interaction or touchpoint that directly led to the conversion. First-Touch Attribution (FTA), as the name suggests, attributes full credit for a conversion to the first interaction or touchpoint that introduced the customer to the brand or campaign. Uniform Attribution (UNI) is an attribution model in advertising measurement that evenly distributes credit for conversions across all touchpoints in a customer’s journey, providing equal weight to each interaction.
Table 5 shows an example of different attribution logics, given the joined Table 4. Since only ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ is relevant to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{2},1}$, both LTA, FTA, and UNI will assign a weight of 1 to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ (similarly for ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{1},1}$). In the case of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{2},2}$, where both ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ and ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},2}$ are relevant, LTA attributes a weight of 1 to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},2}$ as it occurs last. FTA assigns a weight of 1 to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ since it happens first. UNI, on the other hand, evenly distributes the weight, assigning 0.5 to both ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ and ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},2}$.
Because LTA is the most commonly used attribution logic (choi2020online, ), we will use LTA for illustration throughout this paper. However, the techniques discussed can also be applied to other MTA logics, as demonstrated in the experiments (Section 6.3). More sophisticated algorithms, such as incremental value heuristics and Shapley value methods (singal2019shapley, ), will be considered for future work. These algorithms do not focus on a single user’s data but instead consider correlations across the dataset, making their underlying models significantly different from our current settings.
Table 5. Attribution on Day 1 in Table 4
Conversion ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{1},1}$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{2},1}$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{2},2}$
Pub-ID P-1 P-2 P-1 P-2 P-1 P-2
LTA 1 0 1 0 0 1
FTA 1 0 1 0 1 0
UNI 1 0 1 0 0.5 0.5
2.1.2. Measuring Attributed Conversions
With LTA, each conversion is attributed to the impression that occurs last. In order to effectively evaluate the advertisement’s performance across multiple publishers, it is crucial to aggregate the total attribution weights, i.e., total conversion, for each publisher. A high cumulative sum of weights for a publisher signifies that advertisements on that platform have a substantial likelihood of resulting in product conversions. This metric becomes instrumental in assessing the effectiveness of ad campaigns and optimizing advertising strategies.
For a specific publisher and conversion event ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}$, the attribution process operates as follows: If ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}$ is attributed to an impression ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}$, and ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}$ originates from the same publisher (${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}$ occurs within the publisher’s domain), the publisher is assigned an attribution weight of $w=1$. Conversely, if ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}$ is associated with a different publisher, the attribution weight assigned to the original publisher is $w=0$. Then we will compute, for all conversions, the total attribution weights received by each publisher. A high sum of attribution weights for a publisher signifies that advertisements on that platform are more likely to be profitable and convert into actual conversions.
2.1.3. Streaming Measurement
Table 6. Total Conversion via LTA
Pub-ID Day 1 Day 1 to 2 Day 1 to 3 $\cdots$
P-1 $\mathbf{x}_{\text{P1}}[1]$ $\mathbf{x}_{\text{P1}}[1]+\mathbf{x}_{\text{P1}}[2]$ $\sum_{i=1}^{3}\mathbf{x}_{\text{P1}}[i]$cdots
P-2 $\mathbf{x}_{\text{P2}}[1]$ $\mathbf{x}_{\text{P2}}[1]+\mathbf{x}_{\text{P2}}[2]$ $\sum_{i=1}^{3}\mathbf{x}_{\text{P2}}[i]$cdots
In practical scenarios, it is crucial to implement a continuous approach to releasing the sum of attribution weights. This allows for ongoing monitoring of trends over time, enabling the assessment of changes in advertising performance. Table 6 illustrates the cumulative attribution weight, i.e., total conversion, of Nike across two major publishers, Publisher 1 and Publisher 2. This dataset enables the advertiser, Nike, to assess the performance of their advertisements, a crucial factor directly influencing profitability. The continuous online release of attribution data facilitates real-time insights, thereby empowering businesses to make timely and informed decisions that can significantly impact their bottom line.
Table 6 is an example of a private dataset requiring protection. There are $n$ days in total, and we represent the data vector as $\mathbf{x}\in\mathbb{R}^{n}$, where $\mathbf{x}[i]$ represents the sum of attribution weights for a specific publisher on day $i$ ($1\leq i\leq n$). In particular, the sum of attribution weights is equal to the number of conversions if using Last Touch Attribution or First Time Attribution.
For each publisher, we aim to calculate the sum of attribution weights from day 1 to day $i$, offering insights into advertisement performance over a specific timeframe. In mathematical terms, our goal is to release the prefix-sum of $\mathbf{x}$, i.e., $(\mathbf{x}[1],\mathbf{x}[1]+\mathbf{x}[2],\cdots,\sum_{i=1}^{n}\mathbf{x}[i])$. Our primary concern is to ensure that the released information will be privacy-preserving. Specifically, we require that the released information satisfies differential privacy, which will be discussed next in Section 2.2. In Section 4, we will introduce the methodology for achieving this differential privacy-preserving release of the private dataset.
2.2. Differential Privacy
Differential Privacy (dwork06Calibrating, ; dworkKMM06:ourdata, ; zcdp, ; renyidp, ; fdp, ) establishes a set of constraints governing the operation of a mechanism ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$. It has emerged as the de facto standard for safeguarding confidentiality when generating publicly accessible data products. Its adoption extends to a growing number of real-world applications, with notable examples including the U.S. Census Bureau (ashwin08:map, ; onthemap, ; tdahdsr, ), Uber (elasticsensitivity, ; chorus, ), Apple (appledpscale, ), Facebook (fburlshares, ; opacus, ), Microsoft (DingKY17, ), and Google (rappor, ; tensorflowprivacy, ; googlesql, ). Differential privacy offers a robust guarantee of plausible deniability by constraining an adversary’s ability to ascertain whether a specific individual’s data was included in the dataset or not.
One of the key features of differential privacy is its robust privacy guarantee, which remains effective regardless of an adversary’s knowledge or actions when attempting to attack the data. In this context, an adversary refers to any person or entity trying to extract sensitive information about individuals from the results of a data analysis. This privacy guarantee is maintained even if the adversary possesses unlimited computing power and has full knowledge of the algorithm and system used for data collection and analysis.
A fundamental concept in differential privacy is the notion of a neighboring dataset. Let $\mathcal{N}$ represent the set of all possible pairs of neighboring datasets, denoted as $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})$. We define the vector $\mathbf{d}$ as the difference between these neighboring datasets, expressed as $\mathbf{d}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}-\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$. The diffset of $\mathcal{N}$ is defined as the collection of all possible $\mathbf{d}$ values, which we denote as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}}}=\{\mathbf{d}:\mathbf{d}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}-\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}},\forall({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}}}\}$. The most prevalent variant of differential privacy is given below.
Definition 2.1 (Differential Privacy (dworkKMM06:ourdata, )).
Given privacy parameters $\epsilon>0$ and $\delta\in(0,1)$, and the set of all pairs of neighboring datasets ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}}}$, a randomized algorithm ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$ satisfies $(\epsilon,\delta)$-differential privacy if for all $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}}}$ and all sets $S$, the following equations hold:
$P({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})\in S)\leq e^{\epsilon}P({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}(\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\in S)+\delta.$
We introduce an alternative notion of differential privacy, zero-concentrated differential privacy for analysis.
Definition 2.2 (Zero-Concentrated Differential Privacy (zCDP) (zcdp, )).
Given a privacy parameter $\rho>0$ and the set of all pairs of neighboring datasets ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}}}$, a mechanism ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$ satisfies $\rho$-zCDP if for all $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}}}$ and for all $\alpha\in(1,\infty)$, it holds that $DIV_{\alpha}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})||{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}(\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}))\leq\rho\alpha$, where $DIV_{\alpha}$ is the $\alpha$-Renyi divergence between distributions ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})$ and ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}(\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})$.
Theorem 2.3 (zCDP to DP Conversion (canonne2020discrete, )).
If mechanism ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$ satisfies $\rho$-zCDP, then it satisfies $(\epsilon,\delta)$-DP for any $\epsilon>0$ and
$\delta=\min_{\alpha>1}\frac{\exp((\alpha-1)(\alpha\rho-\epsilon))}{\alpha-1}(1-\frac{1}{\alpha})^{\alpha}.$
In this paper, our primary emphasis is on zCDP due to its implementation simplicity. Nevertheless, it’s important to highlight that our proposed method seamlessly extends to other variants of differential privacy. To analyze the privacy of a given mechanism, an important quantity is sensitivity, which measures the maximum change in function value on neighboring datasets.
Definition 2.4.
The $L_{2}$ sensitivity of a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is given by,
$\Delta_{2}(f)=\max_{({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}}}}\|f({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})-f(\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\|_{2},$
where $\|\mathbf{v}\|_{2}$ is the L2 norm of the vector $\mathbf{v}$.
Theorem 2.5 (Gaussian Mechanism of zCDP (zcdp, )).
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ be a function of the dataset ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, the Gaussian mechanism ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$ adds i.i.d. Gaussian noise to $f({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})$ with scale parameter $\sigma^{2}$,
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})=f({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})+N(0,\sigma^{2}\mathbf{I}_{n}),$
where $\mathbf{I}_{n}$ is the $n\times n$ identity matrix. The Gaussian mechanism satisfies $\frac{\Delta_{2}(f)^{2}}{2\sigma^{2}}$-zCDP.
A classical approach is to apply this Gaussian mechanism to the vector of measurement results. However, we later demonstrate that such an i.i.d. distribution is not optimal.
2.2.1. Neighboring Dataset
We first introduce the event-level adjacency ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{Event}}}}$. ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ and $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$ are called event-level neighboring datasets, denoted as $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{Event}}}}$, if $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$ can be obtained from ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ by substituting a single record in ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$. In this context, the diffset of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{Event}}}}$ is defined as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{Event}}}}=\{\mathbf{d}:\|\mathbf{d}\|_{1}\leq 2\}$. This is because when a single record is substituted, $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$ can deviate from ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ by 2 entries, with each entry differing by 1, resulting in $\|\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}\|_{1}\leq 2$.
However, it’s important to recognize that users may not be content with protecting only individual actions taken on the internet. Instead, they often seek comprehensive protection for all of their personal information. Consequently, there arises a need for user-level neighboring datasets. To move from event-level neighboring datasets to user-level neighboring datasets, we replace the concept of "the change of one record" with "the change of all records associated with one user." However, it’s important to note that since a user can potentially contribute an infinite number of records, and differential privacy aims to provide worst-case guarantees, introducing excessive noise can significantly impact utility performance. To address this challenge, a common practice is to constrain user contributions, as discussed in prior works such as (dong2023continual, ; googlesql, ; liu2023algorithms, ). This constraint limits each user to having at most $k$ records in the dataset. In cases where a user exceeds this limit and contributes more than $k$ records, only the first $k$ records are retained, while the surplus records are discarded.
The user-level neighboring dataset with a global contribution constraint of at most $GS$, denoted as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{Global}}}}$, defines a pair of datasets $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})$ as belonging to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{Global}}}}$ if $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$ can be obtained from ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ by substituting all records (up to a maximum of $GS$ records) of one user from ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$. Because each user can contribute a maximum of $GS$ records, the diffset of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{Global}}}}$ can be characterized as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{Global}}}}=\{\mathbf{d}:\|\mathbf{d}\|_{1}\leq 2GS\}$. This is because when $GS$ records are substituted, in the extreme case, $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$ can deviate from ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ by 2 entries, with each entry differing by $GS$, resulting in $\|\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}\|_{1}\leq 2GS$.
2.2.2. Differential Privacy Properties
Theorem 2.6 (Post-processing (zcdp, )).
Let ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$ be a mechanism, $f$ be an arbitrary function that takes the output of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$ as the input. If the mechanism ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$ satisfies $\rho$-zCDP, then the post-processing $f({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}})$ satisfies $\rho$-zCDP.
Theorem 2.7 (Sequential composition (zcdp, )).
Let ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}_{1}$, $\cdots$, ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}_{k}$ be a sequence of mechanisms. If the mechanism ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}_{i}$ in the sequence satisfies $\rho_{i}$-zCDP, then the composition of the sequence ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}=({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}_{1},\cdots,{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}_{k})$ satisfies $(\sum_{i=1}^{k}\rho_{i})$-zCDP.
Theorem 2.8 (Group privacy for zCDP (zcdp, )).
Let ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$ be a mechanism that satisfies $\rho$-zCDP for any event-level neighboring datasets $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{Event}}}}$. Then for any user-level neighboring datasets with a global contribution constraint of at most $k$, $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{Global}}}}$, ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$ satisfies $k^{2}\rho$-zCDP.
Group privacy facilitates the transition from event-level differential privacy to user-level differential privacy. Nevertheless, in the context of online advertising measurement systems, applying group privacy overlooks the likelihood that a user’s contributions are more evenly distributed across each day, rather than being concentrated in a single day. By harnessing the patterns of user contributions across days, it becomes feasible to enhance utility. This is the motivation behind our proposal for a user-level differentially private advertising measurement system with bounded per-day contributions, affording greater control on a day-to-day basis.
3. Related Work
A lack of formal privacy guarantee in industry practice: To address growing concerns over consumer privacy, many advertising networks and platforms have introduced a range of privacy-preserving advertising measurement tools. Notable among these initiatives are Apple’s Private Click Measurement (PCM) framework and Google’s Attribution Reporting API as part of the Privacy Sandbox. PCM mostly relies on reducing data granularity—for example, by allowing only limited user identifiers or implementing delayed reporting—to decrease the amount of information websites can gather about user activities on other sites. Similarly, the Attribution Reporting API delivers event-level reports that share many of PCM’s techniques, such as delayed reporting and coarse data granularity. In addition, the Attribution Reporting API also generates summary reports, or aggregated reports, by utilizing encryption and a trusted execution environment (TEE). Neither PCM nor the Attribution Reporting API can provide any formal privacy guarantee for the advertising measurement output with a mathematically rigorous definition. It is unclear whether such approaches are indeed useful in safeguarding user privacy since any aggregated report could potentially be subject to privacy risks in a way similar to the US census data attack in a series of works by Dwork and others (dwork2017exposed, ; dick2023confidence, ; keller2023database, ). Recently, Google (delaney2024differentially, ) discussed releasing differentially private conversions in a one-shot setting, whereas our focus is on releasing data in an online setting. Alistair (tholoniat2024alistair, ) scheduled the on-device privacy budget efficiently to achieve individual differential privacy (IDP), with a special focus on secure multiparty computation (MPC) and TEE. Additionally, (denison2022private, ) used differentially private stochastic gradient descent to train machine learning models.
Secure computation is not sufficient: Ads industry has recently proposed secure multi-party computation (MPC) based solutions. These solutions ensure the input data encrypted during advertising measurement computations. Examples include PSI-Sum (ion2020deploying, ) by Google and Private Join and Compute (meta, ) by Meta, as well as the Interoperable Private Attribution (IPA) (case2023interoperable, ), a collaborative effort by Mozilla and Meta. Despite the encryption of input data, the reports—output from the MPC computation—are plaintext results without formal privacy guarantees. Recent studies have shown that in the case of advertising measurement, membership inference attacks can exploit this vulnerability, leading to the leakage of user privacy and business confidentiality (guo2022birds, ; Bo, ).
Ensuring privacy guarantees for online reporting poses challenges: An advertising campaign can extend even months. Applying differential privacy to the cumulative ads measurement results ensure the privacy of the final data. However, it is also crucial to report daily measurement results online. Simply applying a differential privacy mechanism by adding independent noise for each day severely degrades measurement accuracy and usability, as demonstrated in the experiments section. One possible research direction to improve the accuracy without sacrificing privacy is to use of tree-based mechanisms to address the online release problem under event-level differential privacy. This research direction was initially introduced by Chan and Dwork (chan2011private, ; dwork2010differential, ). Subsequent studies by Denisov et al. (denisov2022improved, ; henzinger2022constant, ; henzinger2023almost, ) leveraged the matrix mechanism (LMHMR15, ; YZWXYH12, ; li2012adaptive, ; yuan2015optimizing, ; mckenna2018optimizing, ; xiao2020optimizing, ; mckenna2021hdmm, ; xiao2023answering, ) to enhance the accuracy of these tree-based approaches by improving the constant factors in their error rates. (fichtenberger2021differentially, ) studies the continual release problem for graphs.
User level privacy guarantee is required. Two definitions of differential privacy exist in the advertising setting: event-level and user-level. The former safeguards individual events, while the latter conceals all events associated with a particular user across the entire ads measurement campaign. For instance, event-level privacy protects a single user’s click or purchase activity (i.e., conversion), whereas user-level privacy shields all user clicks or conversions. The chosen privacy level impacts the degree of perturbation applied. In this paper, we focus on studying the challenging and more stringent setting of user privacy. Choquette-Choo et al. introduced the matrix mechanism as a solution for user-level differential privacy in their work (choquette2022multi, ; choquette2024amplified, ). Their primary application is differentially private stochastic gradient descent. Nevertheless, it’s worth noting that these methods necessitate a minimum time gap of at least $b$ days between two records belonging to the same user. This requirement presents a challenge in the context of online advertising measurement systems, where a user can potentially have multiple conversions within a single day. Investigating user-level differential privacy in the context of continual observation, (dong2023continual, ; zhang2023differentially, ; george2022continual, ) used tree-based mechanisms to tackle the problem. Although effective for addressing the general streaming problem, these approaches do not leverage the specific structures inherent in advertising measurement systems. For example, in a general streaming process, a single user may contribute most of the records, leading to a large sensitivity (cropping the data related to this user will lead to a high bias). However, in an advertising system involving millions of users, it’s improbable for any individual user to make a substantial contribution to the dataset. In reality, most users generate only a small number of conversions. The removal of data associated with a few users is unlikely to have a significant impact on overall performance.
4. Differentially Private Measurement Design
We will begin by addressing the scenario involving a single publisher. In cases where multiple publishers are involved, we show how to extend the mechanism in Section 4.6. Given a streaming dataset ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ on a publisher for a specific advertiser, where ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[i]$ represents the number of conversions on day $i$, we are interested in releasing a differentially private prefix-sum. For example, given a streaming dataset ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}=({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[1],{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[2],{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[3],{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[4])$, we want to release $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[1],{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[1]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[2],{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[1]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[2]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[3],{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[1]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[2]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[3]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[4])$, which can be represented as a result of matrix-vector product $\mathbf{W}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, where $\mathbf{W}=\left[\begin{smallmatrix}1&0&0&0\\ 1&1&0&0\\ 1&1&1&0\\ 1&1&1&1\end{smallmatrix}\right]$ is called the prefix-sum matrix.
Definition 4.1 (Prefix-sum Matrix).
For a given domain size $n$, a matrix $\mathbf{W}\in\mathbb{R}^{n\times n}$ is called prefix-sum matrix if
$\mathbf{W}[i,j]=\begin{cases}1,&1\leq i\leq j\leq n\\ 0,&\text{otherwise}\end{cases}.$
The next step is to design constraints specific to advertising measurement systems, aimed at preventing improbable scenarios from occurring. This targeted approach enhances accuracy by ensuring that the system operates within realistic bounds, reflecting the unique characteristics of advertising data.
Recalling from Section 2.2.1, the definition of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{Global}}}}$ imposes a global limit on user contributions, overlooking the structures inherent in advertising measurement systems. Within these systems, users with particular interests typically exhibit consistent interactions with ads over time. For instance, sudden bursts of interest followed by complete disengagement are uncommon and may even signal malicious behavior. Instead, it is more typical for a user’s interest in a product to persist, resulting in conversions across multiple days. To capture this behavior effectively, we introduce the concept of bounded per-day contributions.
Definition 4.2 (User-Level Neighboring Dataset with Bounded Per-day Contributions).
Given the constraint that each user can contribute at most $r_{i}$ conversions on day $i$ ($1\leq i\leq n$), the user-level neighboring dataset with bounded per-day contributions, denoted as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$, defines a pair of datasets $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})$ as belonging to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$ if $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$ can be obtained from ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ by substituting all records (with a maximum of $r_{i}$ records on day $i$) of one user from ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$. The diffset of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$ can be expressed as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}=\{\mathbf{d}\in\mathbb{R}^{n}:-r_{i}\leq\mathbf{d}_{i}\leq r_{i},1\leq i\leq n\}$.
4.1. Matrix Mechanism
The matrix mechanism (LMHMR15, ; mckenna2021hdmm, ; xiao2024optimal, ) is a well-established technique for providing differentially private answers to a set of queries. Recently, it has been applied to differentially private online streaming processes (choquette2022multi, ; choquette2024amplified, ), with a particular emphasis on differentially private stochastic gradient descent (DPSGD) (abadi2016deep, ). It factorizes the matrix $\mathbf{W}$ into a strategy matrix $\mathbf{B}$ and a linear transformation matrix $\mathbf{L}$ such that $\mathbf{W}=\mathbf{L}\mathbf{B}$. It first adds independent noise to $\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, then uses $\mathbf{L}$ to transform the noisy answers from $\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ to get noisy answers to $\mathbf{W}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, specifically,
(1) ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})=\mathbf{L}(\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{z})=\mathbf{W}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{L}\mathbf{z}.$
where $\mathbf{z}$ is drawn from a Gaussian distribution $N(0,\mathbf{I}_{n})$, here the covariance matrix $\mathbf{I}_{n}$ is the identity matrix of size $n\times n$.
We aim to minimize the total variance of noise added, for the mechanism defined in Eq. (1). Because ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})=\mathbf{W}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{L}\mathbf{z}$, $\mathbf{z}\sim N(0,\mathbf{I}_{n})$, the total variance of the noise is $\operatorname{trace}(\mathbf{L}\mathbf{L}^{T})$. The matrix mechanism (LMHMR15, ) calculates the matrix $\mathbf{L}$ as $\mathbf{L}=\mathbf{W}\mathbf{B}^{\dagger}$ , where $\mathbf{B}^{\dagger}$ is the Moore–Penrose inverse of $\mathbf{B}$. Here it’s assumed that $\mathbf{B}$ has linearly independent columns so that $\mathbf{B}^{T}\mathbf{B}$ is invertible. Therefore, the total variance can be represented as,
$\operatorname{trace}(\mathbf{L}\mathbf{L}^{T})$ $=\operatorname{trace}(\mathbf{W}\mathbf{B}^{\dagger}(\mathbf{W}\mathbf{B}^{\dagger})^{T})=\operatorname{trace}(\mathbf{W}^{T}\mathbf{W}(\mathbf{B}^{T}\mathbf{B})^{-1}).$
We next analyze the sensitivity of the Mechanism 1. Let the function $f({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})=\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, from Theorem 2.5, to satisfy $\rho$-zCDP, it’s required that the sensitivity is bounded $\Delta_{2}(f)\leq\sqrt{2\rho}$. From the definition of sensitivity in Definition 2.4, and the definition of neighboring dataset in Definition 4.2, the sensitivity of Mechanism 1 is
$\Delta_{2}(f)=\max_{({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}}\|\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}-\mathbf{B}\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}\|_{2}=\max_{\mathbf{d}\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}}\|\mathbf{B}\mathbf{d}\|_{2}\leq\sqrt{2\rho}.$
Notice that $\|\mathbf{B}\mathbf{d}\|_{2}$ is a convex and continuous function with respect to $\mathbf{d}$, and the constraint set ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}$ is convex and compact with respect to $\mathbf{d}$. So according to Bauer’s maximum principle (kruvzik2000bauer, ), the function $\|\mathbf{B}\mathbf{d}\|_{2}$ attains its maximum at some extreme point of the set ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}$. Recall that ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}=\{\mathbf{d}\in\mathbb{R}^{n}:-r_{i}\leq\mathbf{d}_{i}\leq r_{i},1\leq i\leq n\}$, so the set of extreme points of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}$ is
(2) ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}_{ex}}}=\{\mathbf{d}\in\mathbb{R}^{n}:d_{i}\in\{-r_{i},r_{i}\},1\leq i\leq n\}.$
Then the constraint that the sensitivity of Mechanism 1 is at most $\sqrt{2\rho}$ can be written as,
$\max_{\mathbf{d}\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}}\mathbf{d}^{T}\mathbf{B}^{T}\mathbf{B}\mathbf{d}=\max_{\mathbf{d}\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}_{ex}}}}\mathbf{d}^{T}\mathbf{B}^{T}\mathbf{B}\mathbf{d}\leq 2\rho.$
Using the trick in (yuan2015optimizing, ), we change the variable $\mathbf{P}=\mathbf{B}^{T}\mathbf{B}$, then the matrix mechanism problem becomes,
(3) $\min_{\mathbf{P}\succ\mathbf{0}}$ $\operatorname{trace}\left(\mathbf{W}^{\top}\mathbf{W}\mathbf{P}^{-1}\right)$
s.t. $\max_{\mathbf{d}\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}_{ex}}}}\mathbf{d}^{\top}\mathbf{P}\mathbf{d}\leq 2\rho$
The expression $\mathbf{P}\succ\mathbf{0}$ indicates that the matrix $\mathbf{P}$ is positive definite.
4.2. Existing Methods Don’t Scale
When $n$ is relatively small, it’s possible to enumerate the set ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}_{ex}}}$, and the paper (choquette2022multi, ) gives an algorithm to solve the Problem (3). We use the code provided in (choquette2022multi, ) to get the optimal solution $\mathbf{P}^{\star}$ of Problem (3) when $n$ is small, in the hope that we can find some structures in the optimal solution. Note that there are $2^{n}$ vectors in ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}_{ex}}}$, when $n\geq 16$, the code runs out of memory on a Google Colab server with a T4 GPU. It is necessary to get an efficient method to address the scalability issue.
To find a structure in the optimal solution $\mathbf{P}^{\star}$ to Problem (3), we numerically calculate the $\mathbf{P}^{\star}$ using the solver in (choquette2022multi, ) under its default parameters. For each n, we randomly generate the per-day constraint $r_{i}\in[1,10]$ and set $\rho=\frac{1}{2}$. Surprisingly, we find that $\mathbf{P}^{\star}$ tends to be a diagonal matrix. Table 7 shows the absolute value of the off-diagonal elements in $\mathbf{P}^{\star}$, the maximum absolute value of $\mathbf{P}^{\star}$ is close to 0. Under the assumption that the optimal solution $\mathbf{P}^{\star}$ is a diagonal matrix, Problem (3) can be reduced to a one-dimensional optimization problem, which can be efficiently solved when $n$ is large.
Table 7. Check the Absolute Value of the Off-Diagonal Elements.
$n$ $\max_{i\neq j}|\mathbf{P}^{\star}[i,j]|$ $\min_{i\neq j}|\mathbf{P}^{\star}[i,j]|$ $\operatorname{mean}_{i\neq j}|\mathbf{P}^{\star}[i,j]|$
2 $6.67*10^{-5}$ $6.67*10^{-5}$ $6.67*10^{-5}$
4 $1.59*10^{-4}$ $1.14*10^{-5}$ $5.38*10^{-5}$
6 $7.26*10^{-5}$ $8.03*10^{-7}$ $1.82*10^{-5}$
8 $4.81*10^{-5}$ $4.70*10^{-8}$ $5.58*10^{-6}$
10 $4.39*10^{-5}$ $1.69*10^{-8}$ $3.03*10^{-6}$
12 $4.36*10^{-5}$ $9.20*10^{-9}$ $2.14*10^{-6}$
14 $4.48*10^{-5}$ $2.79*10^{-10}$ $1.71*10^{-6}$
4.3. Mechanism Design
Following the conjecture that the optimal solution to Problem 3 is a diagonal matrix, we incorporate independent noises into a scaled version of the data vector ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$. Suppose the length of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ is $n$, let the matrix representation of the scale weight be $\mathbf{B}=diag(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}}\cdots,\frac{1}{\sigma_{n}})$, here $\frac{1}{\sigma_{i}}>0$ is the scale for ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[i]$. Let $\mathbf{L}=\mathbf{W}\mathbf{B}^{-1}$, then $\mathbf{W}=\mathbf{L}\mathbf{B}$.
Recall from Mechanism 1, we first add independent noise to $\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, then use $\mathbf{L}$ to transform the noisy answers from $\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ to get noisy answers to $\mathbf{W}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, specifically,
(4) ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})$ $=\mathbf{L}(\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{z})=\mathbf{L}(diag(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}}\cdots,\frac{1}{\sigma_{n}}){{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{z})$
$=\mathbf{W}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{L}\mathbf{z},$
where $\mathbf{z}$ is drawn from a Gaussian distribution $N(0,\mathbf{I}_{n})$, here the covariance matrix $\mathbf{I}_{n}$ is the identity matrix of size $n\times n$.
The privacy budget is determined by the sensitivity of the function $f({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})=\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, or more precisely, how much the outputs of this mapping can deviate (in terms of $l_{2}$ norm) when transitioning from the input stream ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ to a neighboring stream $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$, under the definition of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$. The following theorem shows how to calculate the sensitivity of the function $f$ with details provided in Appendix 8.1.
Theorem 4.3 (Sesitivity of Ads-BPC).
The sensitivity of the mechanism 4 under the definition of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$ is
$\Delta_{2}(f)=$ $\max_{({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}}\|\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}-\mathbf{B}\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}\|_{2}=\max_{\mathbf{d}\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}}\|\mathbf{B}\mathbf{d}\|_{2}=\sqrt{\frac{r_{1}^{2}}{\sigma_{1}^{2}}+\cdots+\frac{r_{n}^{2}}{\sigma_{n}^{2}}}.$
4.4. Ads-BPC Algorithm
We aim to minimize the total variance of noise added, for the mechanism defined in Eq. 4. Because ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})=\mathbf{W}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{L}\mathbf{z}$, $\mathbf{z}\sim N(0,\mathbf{I}_{n})$, the total variance of the noise is $\operatorname{trace}(\mathbf{L}\mathbf{L}^{T})$. Recall that we calculate the matrix $\mathbf{L}$ as $\mathbf{L}=\mathbf{W}\mathbf{B}^{-1}$, where $\mathbf{B}$ is a diagonal matrix. Therefore, the total variance can be represented as,
$\operatorname{trace}(\mathbf{L}\mathbf{L}^{T})$ $=\operatorname{trace}(\mathbf{W}\mathbf{B}^{-1}(\mathbf{W}\mathbf{B}^{-1})^{T})$
$=\operatorname{trace}(\mathbf{W}^{T}\mathbf{W}(\mathbf{B}^{T}\mathbf{B})^{-1})$
$=n\sigma_{1}^{2}+(n-1)\sigma_{2}^{2}+\cdots+\sigma_{n}^{2}.$
The last equation holds because $(\mathbf{W}^{T}\mathbf{W})[i,i]=n-i+1,1\leq i\leq n$.
From Theorem 2.5, to satisfy $\rho$-zCDP, it’s required that the sensitivity is bounded $\Delta_{2}(f)\leq\sqrt{2\rho}$. Because we want to achieve $\rho$-zCDP while minimizing the total variance, the problem becomes,
(5) $\min_{\sigma_{1}^{2},\cdots,\sigma_{n}^{2}}\quad$ ${n}\sigma_{1}^{2}+(n-1)\sigma_{2}^{2}+\cdots+\sigma_{n}^{2}$
$s.t.\quad$ $\frac{r_{1}^{2}}{\sigma_{1}^{2}}+\frac{r_{2}^{2}}{\sigma_{2}^{2}}+\cdots+\frac{r_{n}^{2}}{\sigma_{n}^{2}}\leq 2\rho.$
The following theorem shows that matrix $\mathbf{B}$ derived above satisfies $\rho$-zCDP. The proof is shown in Appendix 8.2.
Theorem 4.4.
Let $\mathbf{W}\in\mathbb{R}^{n\times n}$ be the prefix-sum matrix 4.1. Let $\sigma_{1},\sigma_{2},$ $\cdots\sigma_{n}$ be the solution to problem 5, $\mathbf{B}=diag(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}}\cdots,\frac{1}{\sigma_{n}})$ and $\mathbf{L}=\mathbf{W}\mathbf{B}^{-1}$. Let $\mathbf{z}\sim N(0,\mathbf{I}_{n})$, then ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})$ = $\mathbf{L}(\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{z})=\mathbf{W}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{L}\mathbf{z}$ satisfies $\rho$-zCDP.
Problem (5) needs to be solved efficiently for the online service. In Appendix 8.3, we obtain the following results:
Theorem 4.5.
Given the optimization problem in Problem (5), the optimal objective function value is $J=\left(\sum_{i=1}^{n}\sqrt{(n-i+1)r^{2}_{i}}\right)^{2}/(2\rho)$ with the optimal solution being $\sigma^{2}_{i}=\sqrt{(Jr^{2}_{i})/(2\rho(n-i+1))}$.
Note that the optimal noise distribution is non-i.i.d. distribution. Before introducing the noise addition algorithm, we outline our data truncation process, which limits each user’s contributions per day. For a fixed publisher, the truncation process operates at each timestamp $k$, tracking conversions ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{k}$ and their corresponding weights $w_{k}$. These weights indicate whether the conversion is attributed to impressions belonging to the given publisher. Specifically, $w_{k}=1$ if ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{k}$ is attributed to impressions from the given publisher, and $w_{k}=0$ if ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{k}$ is attributed to impressions from other publishers.
We maintain a dictionary, user_contribution, to aggregate the sum of attribution weights for each user. If a user’s sum of attribution weights exceeds a pre-specified threshold $r$, only the first $r$ sums of attribution weights will be retained in the dataset; any subsequent weights will be truncated. The detailed algorithm is presented in the Appendix 8.6.
Algorithm 1 outlines the process for releasing the user-level differentially private sum of attribution weights for a single publisher each day. In Line 1, we calculate the scale parameter using Theorem 4.5. Subsequently, in Line 1, we introduce standard Gaussian noise to the scaled version of the truncated ad measurements. This step is equivalent to adding noise from a $N(0,\sigma_{i}^{2})$ distribution to the truncated ad measurements. Finally, Line 1 reconstructs the total sum of attribution weights from day 1 to day $i$ and releases this private information.
Input: Number of days $n$, privacy budget $\rho$, the maximum user contributions at day $i$ is $r_{i}$, the truncated ads measurement on day $i$ is $x[i]$.
Output: The noisy ads measurement from day 1 to day $i$.
1 $J=\left(\sum_{i=1}^{n}\sqrt{(n-i+1)r^{2}_{i}}\right)^{2}/(2\rho)$
2 for $i=1\to n$ do
3       $\sigma^{2}_{i}=\sqrt{(Jr^{2}_{i})/(2\rho(n-i+1))}$
4       $y_{i}=\frac{1}{\sigma_{i}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[i]+z_{i}$, $\quad z_{i}\sim N(0,1)$
5       Release $\sum_{t=1}^{i}\sigma_{t}*y_{t}$
6
Algorithm 1 Ads-BPC
4.5. Sanity Check for Proposed Method
Under the same setting of the small-scale numerical experiments in Table 7, we compare the optimal objective function values of the solver in (choquette2022multi, ) and our method. Table 8 shows that our method achieves optimality up to a difference of $10^{-2}$, which may come from numerical error.
4.6. Extension for Multiple Publishers
The previous sections assume that only one publisher is involved. In practice, multiple publishers often run the ad campaign. In this section, we demonstrate that it is straightforward to extend the previous algorithm to accommodate the multi-publisher case.
Suppose there are $k$ publishers, the data vector for publisher $i$ is ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}_{i}\in\mathbb{R}^{n\times 1}$. Let $\mathbf{X}\in\mathbb{R}^{n\times k}$ be the concatenation of these $k$ data vectors ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}_{i}$. Similar to Section 4.3, the goal is to find a private answer to $\mathbf{W}\mathbf{X}$. We factorize the workload matrix $\mathbf{W}=\mathbf{L}\mathbf{B}$, where $\mathbf{B}=diag(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}}\cdots,\frac{1}{\sigma_{n}})$ and $\mathbf{L}=\mathbf{W}\mathbf{B}^{-1}$. Let $\mathbf{Z}\sim MVG_{n,k}(0,\mathbf{I}_{n},\mathbf{I}_{k})$, where $MVG_{n,k}(0,\mathbf{I}_{n},\mathbf{I}_{k})$ is a Matrix Gaussian Distribution (chanyaswad2018mvg, ; yang2021improved, ), then the mechanism is
(6) ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}(\mathbf{X})=\mathbf{L}(\mathbf{B}\mathbf{X}+\mathbf{Z})=\mathbf{W}\mathbf{X}+\mathbf{L}\mathbf{Z}.$
Given the constraint that each user can contribute at most $r_{i}$ records on day $i$ ($1\leq i\leq n$), the user-level neighboring dataset with bounded per-day contributions, denoted as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$, defines a pair of datasets $(\mathbf{X},\tilde{\mathbf{X}})$ as belonging to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$ if $\tilde{\mathbf{X}}$ can be obtained from $\mathbf{X}$ by substituting all records (with a maximum of $r_{i}$ records on day $i$) of one user from $\mathbf{X}$. The diffset of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$ can be expressed as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}=\{\mathbf{D}\in\mathbb{R}^{n\times k}:0\leq\|\mathbf{D}[i,:]\|_{1}\leq r_{i},1\leq i\leq n\}$.
The sensitivity calculation is similar, given that the user contribution for each day is at most $r_{i}$ (across all publishers),
Theorem 4.6 (Sesitivity of Mechanism 6).
The sensitivity of the mechanism 6 under the definition of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$ is
(7) $\Delta_{2}(f)=$ $\max_{(\mathbf{X},\tilde{\mathbf{X}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}}\|\mathbf{B}\mathbf{X}-\mathbf{B}\tilde{\mathbf{X}}\|_{F}=$ $\sqrt{\frac{r_{1}^{2}}{\sigma_{1}^{2}}+\cdots+\frac{r_{n}^{2}}{\sigma_{n}^{2}}}.$
Then similar to Theorem 4.4, we can show that the mechanism 6 also satisfies $\rho$-zCDP,
Theorem 4.7.
Let $\mathbf{W}\in\mathbb{R}^{n\times n}$ be the prefix-sum matrix 4.1. Let $\sigma_{1},\sigma_{2},$ $\cdots\sigma_{n}$ be the solution to problem 5, $\mathbf{B}=diag(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}}\cdots,\frac{1}{\sigma_{n}})$ and $\mathbf{L}=\mathbf{W}\mathbf{B}^{-1}$. Let $\mathbf{Z}\sim MVG_{n,k}(0,\mathbf{I}_{n},\mathbf{I}_{k})$, then ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}(\mathbf{X})$ = $\mathbf{L}(\mathbf{B}\mathbf{X}+\mathbf{Z})=\mathbf{W}\mathbf{X}+\mathbf{L}\mathbf{Z}$ satisfies $\rho$-zCDP.
Table 8. Comparison between the optimal objective function value of different methods.
$n$ Solver Our Method $n$ Solver Our Method
4 950.55 950.56 10 11365.32 11365.32
6 2400.11 2400.12 12 19441.07 19441.07
8 5193.02 5193.03 14 32607.31 32607.32
5. Discussion
In our specific context, it’s important to note that two data streams, denoted as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ and $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$, are considered neighbors when one can be derived from the other through the substitution of one user’s data. Notably, this solution can be derived from the traditional definition of neighboring datasets, where ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ and $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$ are regarded as neighbors when one can be obtained from the other by either adding or removing one user’s data. This transformation arises due to the possibility of utilizing twice the privacy budget: first for removing a user’s data and then for adding another user’s data.
In this paper, we primarily focus on rule-based attribution logics, including LTA, FTA, UNI, and other widely used multi-touch attribution methods. These methods assign attribution weights based on each user’s data. Methods that incorporate data across multiple users, such as incremental value heuristics and Shapley value (SV) methods, are left for future work.
It’s essential to recognize that our approach assumes the availability of upper bounds for user contributions, denoted as $r_{i}$. If a user surpasses the limit $r_{i}$ with records on day $i$, only the initial $r_{i}$ records will be retained in the dataset, while the excess will be truncated. Determining the appropriate values for $r_{i}$ entails navigating the bias-variance trade-off. Opting for a small $r_{i}$ leads to high bias, whereas a large $r_{i}$ introduces increased variance. Alternatively, a portion of the privacy budget could be allocated for examining the dataset and approximating the values of the upper bounds, $r_{i}$. However, we defer this task to future research, as the techniques involved in this estimation process are orthogonal to our current method.
Furthermore, it’s worth noting that our method offers robustness against malicious inputs exceeding the specified cap for individual contributions. This safeguard ensures that a single report cannot compromise the integrity of a query’s outcome by introducing unbounded trigger values.
6. Experiments
Throughout the experiments, we assume the length of the data vector is $n=len({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})$, that is, we need to report the accumulated attribution weights for $n$ days. For each dataset, we assume the global sensitivity is given as $GS$. In the case where the global sensitivity is unknown, we can set a reasonably large $GS$ as a threshold. Any user with more than $GS$ records in the dataset will only have the first $GS$ records retained, and any additional records will be removed.
On each day, we collect the relative error, then we report their median and 90% maximum (90th quantile). We will also report the root mean squared error (RMSE). For experiment with more than 1 publisher, we report the median error across all publishers. We repeat our experiments 10 times and report the average results. All experiments are conducted on a desktop with an 8-core 16-thread 3.6GHz Intel i9-9900K CPU and 32 GB memory.
6.1. Baseline Methods
We compare our algorithm (named Ads-BPC) with other baselines. Unless otherwise specified, we allocate a total privacy budget of $\rho_{total}=1$. We briefly introduce each baseline below and provide detailed descriptions in Appendix 8.7.
IPA (case2023interoperable, ): The Interoperable Private Attribution (IPA) protocol leverages a combination of multi-party computation and differential privacy to facilitate the processing of individuals’ data while ensuring that only aggregated measurements are disclosed. In this paper, our primary focus is on comparing how IPA achieves differential privacy, assuming the presence of a trusted third party that provides the joined table. IPA adds independent noise to the data vector ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, we use the group privacy property 2.8 to make sure it satisfies user-level neighboring datasets with a global contribution constraint of at most $GS$.
Id-BPC (add independent noise with bounded per-day contribution): We set a maximum number of contributions $r$ a user can make in 1 day, if any user contributes more than $r$ records (conversions) in 1 day, only the first $r$ records will be kept in the dataset, other records are cropped. Then we add independent noise to the data vector ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ (with proper scales). Id-BPC is proposed based on IPA to check the influence of bounded per day contributions.
StreamUDP (dong2023continual, ): StreamUDP adds independent noise to a binary tree structure, we use the group privacy property 2.8 to make sure it satisfies user-level neighboring datasets with a global contribution constraint of at most $GS$.
Bin-BPC (binary tree method with bounded per-day contribution): Similar to Id-BPC, we set a maximum number of contributions $r$ a user can make in 1 day. Then we add independent noise to a binary tree structure (with proper scales). Bin-BPC is proposed based on StreamUDP to check the influence of bounded per day contributions.
UMM (user-level matrix mechanism): It’s adapted from the event-level matrix mechanism (henzinger2022constant, ), we use the group privacy property 2.8 to make sure it satisfies user-level neighboring datasets with a global contribution constraint of at most $GS$.
6.2. Evaluation on Synthetic Datasets
In this section, we generate synthetic datasets to test the performance of our algorithm. We assume there are 1 million users, $|U|=10^{6}$. There are 1000 publishers and the privacy budget is $\rho=1$. For each user, we use 3 different distributions to generate the number of total conversions in the dataset (the distributions are similar to the settings in (dong2023continual, )). In this section, we assume last-touch attribution (LTA) is used. For example, if the distribution generates a number 5 for the user Alice, it means that Alice has 5 conversions in the dataset.
Since it’s improbable for a user to interact with all 1000 publishers, we opt to generate a random number $q$ within the range [1, 100] for each user, denoting the potential publishers they could interact with. Subsequently, we randomly select $q$ publishers from the pool of 1000 as the candidate set. For each conversion made by a user, we then randomly assign it to one of these $q$ publishers. Following this, we shuffle the dataset and evenly divide it into $n$ segments, where the i-th segment represents the conversions received on day i. Table 9 shows an example of the dataset. At each time (we assume each conversion happens in a non-overlapped timestamp), there is one conversion, that’s attributed to one of the publishers.
Time 1 2 3 4 $\cdots$
User Alice Bob Alice Bob $\cdots$
Publisher 1 0 1 0 0 $\cdots$
Publisher 2 1 0 0 0 $\cdots$
$\cdots$ $\cdots$ $\cdots$ $\cdots$ $\cdots$ $\cdots$
Publisher 1000 0 0 0 1 $\cdots$
Table 9. Attribution weight for each publisher
For the algorithms StreamUDP and IPA, the global sensitivity of the dataset is denoted as $GS$. As for Bin-BPC, Id-BPC, and Ads-BPC, we set the threshold for daily contributions as $r$. Selecting an appropriate threshold value $r$ involves a trade-off between bias and variance. A small $r$ leads to the removal of numerous records, introducing substantial bias. Conversely, a large $r$ retains most records in the dataset but introduces significant noise into the results, resulting in high variance. In this paper, we manually determine the threshold $r$, while considering the adaptive selection of $r$ as a potential avenue for future research. The adaptive determination of $r$ is considered a parallel task to our current method, and any suitable technique (e.g., as discussed in (dong2023continual, )) can be applied to enhance our approach. We use $r=\lceil\frac{GS}{n}*2\rceil$ as the threshold, with the hope that most of the users have contributions less than $r$ on each day.
6.2.1. Experiment on Zipf Dataset
In the Zipf dataset, the number of conversions for each user is generated from the distribution: Zipf(3)+10. Here the global sensitivity is GS = 50 (A user has at most 50 conversions for each publisher). The total number of conversions in the dataset is $T=11,361,742$. The user contribution limit for each day is set as $r=\lceil\frac{50}{31}*2\rceil=4$. Table 10 displays the median relative error, 90% maximum relative error, and root mean squared error across different publishers for various algorithms. Ads-BPC stands out for achieving the lowest error among all algorithms. It demonstrates an improvement of approximately 35% over the second-best algorithm, Id-BPC.
Error StreamUDP BinBPC IPA IdBPC UMM AdsBPC
Median Relative Error (%) 1.61 1.20 2.66 0.83 1.70 0.55
90% Max Relative Error (%) 4.95 4.63 6.16 1.94 7.21 1.21
RMSE 126.08 94.72 174.71 54.53 127.27 37.62
Table 10. Error Comparison on Zipf Dataset
Error StreamUDP BinBPC IPA IdBPC UMM AdsBPC
Median Relative Error (%) 1.09 0.68 1.80 0.47 1.15 0.24
90% Max Relative Error (%) 3.35 2.61 4.18 1.10 4.87 0.54
RMSE 379.58 237.45 522.20 138.62 380.86 73.95
Table 11. Error Comparison on Normal Dataset
Error StreamUDP BinBPC IPA IdBPC UMM AdsBPC
Median Relative Error (%) 0.73 0.49 1.21 0.32 0.77 0.15
90% Max Relative Error (%) 2.24 1.97 2.82 0.84 3.27 0.36
RMSE 650.59 625.32 900.86 535.92 651.58 488.21
Table 12. Error Comparison on Uniform Dataset
6.2.2. Normal distribution
The number of conversions for each user is generated from the distribution: Normal(50, 30). Here the global sensitivity is GS = 150. The total number of conversions in the dataset is $T=50,193,976$. The user contribution limit for each day is $r=\lceil\frac{150}{31}*2\rceil=10$. In Table 11, Ads-BPC consistently outperforms other algorithms, showcasing the lowest median relative error, 90% maximum relative error, and root mean squared error. On average, Ads-BPC demonstrates a 48% improvement over Id-BPC, which ranks second in performance.
6.2.3. Uniform distribution
The number of conversions for each user is generated from the distribution: Uniform(1, 256). Here the global sensitivity is $GS=256$, The total number of conversions in the dataset is $T=51,180,247$. The user contribution limit for each day is $r=\lceil\frac{256}{31}*2\rceil=17$. Table 12 illustrates that Ads-BPC consistently outperforms other algorithms across various publishers, showcasing an average improvement of 50% over the second-best algorithm, Id-BPC.
6.3. Experiments on Criteo Dataset
The Criteo Sponsored Search Conversion Log Dataset (tallis2018reacting, ) contains logs obtained from Criteo Predictive Search (CPS). It offers an automated end-to-end solution using sophisticated machine learning techniques to enhance the Google Shopping experience through robust predictive optimization across all aspects of an advertiser’s campaign. Each row in the dataset represents a user action, such as a click, performed on a product-related advertisement. These product advertisements are displayed to users after they express intent through an online search engine. Each row in the dataset includes information about product characteristics (age, brand, gender, price), the timing of the click, user characteristics, and device information. Additionally, the logs contain information about whether these clicks eventually led to a conversion, which means the product was purchased within an $n$-day window, along with the time elapsed between the click and the conversion.
[…]
We use different attribution logic (LTA, FTA, UNI) in the experiment, and the results show that Ads-BPC performs the best regardless of the attribution logic. Table 13 presents the median relative error, 90% maximum relative error, and root mean squared error of different algorithms across various publishers when LTA is used. Notably, the Ads-BPC algorithm attains the lowest error compared to other algorithms, leading to an improvement of approximately 25% over Id-BPC. Table 14 shows the errors of different algorithms when FTA is used. Ads-BPC has the lowest error compared to other algorithms, with an improvement of approximately 30% over Id-BPC. Using UNI attribution logic, Table 15 presents the errors of different algorithms. Ads-BPC improves 33% over Id-BPC.
Error StreamUDP BinBPC IPA IdBPC UMM AdsBPC
Median Relative Error (%) 11.90 7.65 20.63 5.25 13.37 4.02
90% Max Relative Error (%) 38.15 30.14 48.35 12.69 57.37 9.34
RMSE 111.40 71.25 150.08 40.97 112.93 29.88
Table 13. Error Comparison on Criteo Dataset using LTA
Error StreamUDP BinBPC IPA IdBPC UMM AdsBPC
Median Relative Error (%) 7.78 5.02 12.06 3.37 8.05 2.38
90% Max Relative Error (%) 22.89 18.77 29.99 7.53 34.15 5.42
RMSE 112.51 70.42 154.09 40.80 111.85 30.37
Table 14. Error Comparison on Criteo Dataset using FTA
Error StreamUDP BinBPC IPA IdBPC UMM AdsBPC
Median Relative Error (%) 7.94 4.84 12.16 3.42 8.42 2.28
90% Max Relative Error (%) 23.78 18.84 30.05 8.18 36.10 5.41
RMS 111.09 71.31 151.98 41.16 112.39 30.13
Table 15. Error Comparison on Criteo Dataset using UNI
6.4. Experiments on Facebook Sales Conversion Dataset
Facebook Sales Conversion Dataset (fbdata, ) is a dataset used for Facebook ad campaign analysis and sales prediction. Each row of the raw dataset contains information about age, gender, interest, and number of conversions from a unique user group. We treat each unique user group as a single user and distribute the conversions across $n$ days. Key dataset statistics include $\text{GS}=60$, $T=3264$, and $|U|=1143$. The conversions are distributed to $n=31$ days. The user contribution limit for each day is $r=\lceil\frac{60}{31}*2\rceil=4$. The privacy budget is $\rho=1$. Table 16 presents the median relative error, 90% maximum relative error, and root mean squared error of different algorithms across various publishers when LTA is used. Ads-BPC algorithm has an improvement of approximately 50% over the second best Id-BPC.
Error StreamUDP BinBPC IPA IdBPC UMM AdsBPC
Median Relative Error (%) 5.26 4.21 11.36 2.81 8.80 1.42
90% Max Relative Error (%) 18.75 15.46 27.48 6.68 31.99 3.27
RMSE 135.46 91.10 206.01 54.42 166.61 30.19
Table 16. Error Comparison on Facebook Sales Conversion Dataset
6.5. Experiments on Different Parameters
In this section, we demonstrate how different parameters influence the errors. We use the Criteo dataset for illustration. The results for the Facebook Sales Conversion dataset are provided in the appendix.
6.5.1. Experiments on Different Ad Campaign Durations
Figure 1. Comparison of different number of days $n$ on Criteo Dataset.
Figure 2. Comparison of different privacy budgets $\rho$ on Criteo Dataset.
Figure 3. Comparison of different user contributions $r$ on Criteo Dataset.
In this section, we study the performance of different algorithms with respect to the number of days $n$, $n$ is chosen from [7, 15, 31, 63, 127, 255]. The dataset we use is the Criteo dataset with LTA as described in Section 6.3. The privacy budget is set as $\rho=1$. The user contribution limit for each day is $r=\lceil\frac{44}{n}*2\rceil$. Figure 1 illustrates the performance of various algorithms. Ads-BPC outperforms others consistently.
6.5.2. Experiment on Different Privacy Budgets
In this section, we study the performance of different algorithms with respect to the privacy budget $\rho$, $\rho$ is chosen from [0.25, 0.5, 1, 2, 4, 8]. The dataset we use is the Criteo dataset with LTA as described in section 6.3. The number of days is $n=31$. The user contribution limit for each day is $r=\lceil\frac{44}{31}*2\rceil=3$. As depicted in Figure 2, the errors associated with various algorithms decrease as $\rho$ increases. This observed trend can be attributed to the fact that with a larger privacy budget, the amount of noise added to the dataset decreases, leading to lower error rates.
6.5.3. Experiments on Different User Contributions
In this section, we study the performance of different algorithms with respect to the user contribution limits $r$, $r$ is chosen from [1, 2, 3, 4, 5, 6, 8, 10, 12]. The dataset we use is the Criteo dataset with LTA as described in Section 6.3. The number of days is $n=31$. The privacy budget is $\rho=1$. In Figure 3, as the user contribution limit increases, more noise is introduced into the data, resulting in higher variance. Ads-BPC consistently beats other baseline methods.
7. Conclusions
Our work is the first to formulate the real-world problem of MTA advertising measurement systems with real-time reporting of streaming data in advertising campaigns. We introduced Ads-BPC, a novel differentially private solution in this context, which employs a matrix mechanism formulation and further minimizes the total noise power with an analytical solution, leading to non-i.i.d. noise distribution for streaming differentially private measurement results. Extensive experiments confirm that Ads-BPC significantly enhances measurement accuracy by 25% to 50% compared to existing streaming differential privacy mechanisms applied in advertising measurement. This improvement underscores the effectiveness of our approach in achieving superior accuracy while maintaining a formal privacy guarantee, thus advancing the state-of-the-art in privacy-preserving advertising measurement.
Acknowledgments
The authors used ChatGPT4o to revise the text in Section 4 and Section 6 to “Revise the following text, keep the latex syntax, fix typos and make it easy to understand".